Find the solutions of 3x2/3 + 8x1/3 - 3 = 0.(eg.3 with u = x 1/3)
College Algebra ISBN-13: 978-0-618-77629-0
Thanks

Answer by stanbon(57258)

Find the solutions of 3x2/3 + 8x1/3 - 3 = 0
----------
You have a quadratic with variable x^(1/3).
-------------------------
Let x^1/3 = w
Substitute to get:
3w^2 + 8w - 3 = 0
Use the AC Method or "Reverse Foil" to factor the left side:
------------
3w^2+9w-w-3 = 0
3w(w+3)-(w+3) = 0
(w+3)(3w-1) = 0
w = -3 or w = 1/3
------------
Substitute to get:
x^(1/3) = -3 or x^(1/3) = 1/3
x = (-3)^3 or x = (1/3)^3
x = -27 or x = 1/27
========================
Cheers,
Stan H.